# MultiSocial: Multilingual Benchmark of Machine-Generated Text Detection of Social-Media Texts

### v1

@@ Abstract
Recent LLMs are able to generate high-quality multilingual texts, indistinguishable for humans from authentic human-written ones. Research in machine-generated text detection is however mostly focused on the English language and longer texts, such as news articles, scientific papers or student essays. Social-media texts are usually much shorter and often feature informal language, grammatical errors, or distinct linguistic items (e.g., emoticons, hashtags). There is a gap in studying the ability of existing methods in detection of such texts, reflected also in the lack of existing multilingual benchmark datasets. To fill this gap we propose the first multilingual (22 languages) and multi-platform (5 social media platforms) dataset for benchmarking machine-generated text detection in the social-media domain, called MultiSocial [^1] https://anonymous.4open.science/r/multisocial. It contains 472,097 texts, of which about 58k are human-written and approximately the same amount is generated by each of 7 multilingual LLMs. We use this benchmark to compare existing detection methods in zero-shot as well as fine-tuned form. Our results indicate that the fine-tuned detectors have no problem to be trained on social-media texts and that the platform selection for training matters.
@@ 1 Introduction
The most advanced text-generation AI models, called large language models (LLMs), are able to generate high-quality texts in various languages (Qin et al., 2024). Although this presents an opportunity to make a human life and work more efficient, it also presents a threat of being misused, as such generated texts are not easily recognisable by humans (Zellers et al., 2019). This is especially crucial in regard to social-media networks (SMN, such as Facebook, X/Twitter, etc.), where anyone can be a source of a harmful content (without editorial consent) with a potentially wide reach (depending on a network) (Aïmeur et al., 2023). To prevent the LLM misuse (e.g., social engineering, disinformation spreading), a reliable mechanism to detect machine-generated text (MGT) is needed.
@@ 2 Related Work
Benchmark datasets HC3 (Guo et al., 2023) and SAID (Cui et al., 2023) contain texts Chinese and English covering forum-like question-answering domain. HC3 contains only ChatGPT-generated machine texts, SAID contains real bot-generated texts (relying on human annotations to identify them). The downside of SAID is that the specific generation model of individual texts is not known. Otherwise, social-media texts are covered only in rather old TweepFake (Fagni et al., 2021) dataset, which includes English-only tweets and cannot be used to evaluate MGT detection methods on social-media texts generated by the most modern LLMs. Fox8-23 (Yang and Menczer, 2023) dataset also focuses on English, and covers presumably only ChatGPT-generated tweets. Similarly, F3 (Lucas et al., 2023) dataset contains English ChatGPT-generated real and fake news as well as tweets. There are other monolingual works, such as (Temnikova et al., 2023) focused on fake tweets in Bulgarian; however, the crafted dataset is not publicly available.
@@ 3 Dataset
Since there is no dataset of multilingual SMN texts containing human-written texts along with the texts generated by SOTA text-generation LLMs, we have crafted a new MultiSocial benchmark dataset. It contains human-written data from five different social-media platforms (reused from the existing multilingual SMN datasets), namely Telegram, Twitter (X), Gab, Discord, and WhatsApp, including post-like as well as chat-like texts. For each authentic human-written text, the texts generated by 7 SOTA LLMs are included by using three iterations of paraphrasing. Details regarding dataset construction (including selection, pre-processing, generation, and post-processing of texts) are provided in Appendix B.
Dataset consists of 472,097 texts (58k are human-written) split into train and test subsets, of which sample counts are summarized across languages and platforms in Table 2. We have conducted linguistic analysis of the machine-generated texts along with the similarity comparison to human texts, with a manual human check of a balanced subset, and identified small portion (about 1%) of noise in the generated data (e.g., “As an AI model…”), indicating model failure during generation. We have intentionally left such samples in the dataset (clearly marked in the data) for further analysis purpose (as indicated in Appendix B). We however filter-out the identified noise for the purpose of the experiments in this study. Although such a post-processing cannot guarantee 100% removal of noisy data, the obvious failures are cleaned.
@@
The MultiSocial benchmark dataset covers 22 (high- and low-resource) languages (some only in the testing split), 20 of which (18 of Indo-European and 2 of Uralic language families) have been selected based on our research-projects needs. However, we have intentionally included 2 more (Arabic of Semitic and Chinese of Sino-Tibetan family), which are completely linguistically and geographically unrelated, to study cross-lingual characteristics (Figure 1).
The dataset is strongly focused on the Indo-European language family, but contains 4 language families in total. Test split includes all of the train languages, but also 2 Celtic languages, 1 more South Slavic and 1 more West Slavic language. There are 5 writing scripts in both train and test splits of the dataset, where majority of languages uses Latin, but there is also Cyrillic (Russian, Ukrainian, Bulgarian), Arabic, Hanzi (Chinese), and Greek. Nice feature is that the train split includes at least 3 representatives of Germanic, Romance, Slavic-Latin, and Slavic-Cyrillic, which enables various combinations of studies regarding multilingual and cross-lingual characteristics of machine-generated text detection. Furthermore, in Slavic and Romance languages, there are included at least 2 representatives of languages that can be considered high-resource and low-resource.
Although not all languages have enough samples from each of five platforms, specific subsets can be used to study specific characteristics (e.g., cross-platform transferability using English and Spanish languages, cross-lingual transferability using Telegram platform, surprise language and/or surprise platform evaluation not using specific portions of the train data, etc.).
@@ 4 Detection Methods
For statistical category, we have selected the following 5 most-promising methods (excluding perturbation-based and multi-generation methods due to high computing costs): Binoculars (Hans et al., 2024) with Falcon-7B (Almazrouei et al., 2023) as an observer model and Falcon-7B-Instruct as a performer model, Fast-DetectGPT (Bao et al., 2023) with GPT-J-6B (Wang and Komatsuzaki, 2021) as both the reference and sampling models, LLM-Deviation (Wu and Xiang, 2023), DetectLLM-LRR (Su et al., 2023), S5 (Spiegel and Macko, 2024) (multiplying 5 statistical metrics of Likelihood, Entropy, Rank, Log-Rank, and LLM-Deviation), all three of them using GPT-J-6B as a base model (the same as in Fast-DetectGPT).
@@
For fine-tuned category, we have selected 7 multilingual foundational models covering multiple architectures and model sizes: mDeBERTa-v3-base (He et al., 2022) (as the best detector in Macko et al., 2023), XLM-RoBERTa-large (Conneau et al., 2020) (as the best detector in Macko et al., 2023 based on AUC ROC), Mistral-7B (Jiang et al., 2023) (as the best single-model multilingual detector in of SemEval-2024 Task 8 Wang et al., 2024b), Llama-3-8B (AI@Meta, 2024) (as a SOTA multilingual smaller LLM), Aya-101 (Üstün et al., 2024) (as a SOTA representative of multilingual LLMs with encoder-decoder architecture), BLOOMZ-3B (Muennighoff et al., 2022) (as the best pre-trained detector base model), and Falcon-rw-1B (as a smaller version of the best performing model at ALTA 2023 Gagiano and Tian, 2023, since the 7B model version is already covered by better performing Mistral, Spiegel and Macko, 2024).
@@ 5 Experimental Results
Table 3: Benchmark of the selected MGTD methods of statistical, pre-trained, and fine-tuned categories.
@@ 6 Discussion
Selection of social-media platform for fine-tuning matters. There are also significant differences between models trained on single vs multiple platform dataset. For example, on Twitter data, the Discord-trained detectors achieve on average 13% lower AUC ROC than Telegram-trained detectors.
@@ 7 Conclusions
We have created and published a unique multi-platform and massively multilingual dataset, named MultiSocial, to benchmark machine-generated text detection methods on social-media texts. It covers 7 most modern text-generation AI models (of various sizes and architectures), 5 social-media network platforms, and 22 languages of 4 primary language families. We have used this dataset to benchmark 17 carefully selected state-of-the-art detection methods of 3 categories (statistical zero-shot, pre-trained, and fine-tuned) and compare their multi-platform and multi-lingual capabilities (as well as cross-lingual and cross-platform capabilities of fine-tuned detectors). We have discussed the most interesting findings, including that the detection models can be fine-tuned to the machine-generated text detection task using social-media texts (shorter lengths, informal style, emoticons and hashtags) quite well, with the performance comparable to the performance reported in other domains (e.g., news articles).
Due to rather high performance differences in cross-platform evaluation, the further work should be focused to a more-detailed analysis of cross-domain multilingual capability of the state-of-the-art detectors. The proposed MultiSocial dataset can be used for a more detailed multilingual evaluation as well, such as a selection of optimal minimal subset of languages and platforms for training.
@@
Biases. Although the dataset contains a wide variety of languages (22 in total) covering various scripts and language families (see Section 3), the dataset is still biased towards Indo-European languages with 18 out of 22 belonging to this family. The dataset also reflects the topics characteristic for the time and the social media included in the 6 original datasets used as sources of human-written texts, but they should already be rather varied due to sheer volume of the data included (see Appendix B).
Misuse Potential. We work with already publicly available datasets of human-written social media content as well as with publicly available LLMs to generate the texts. In general, the human-written texts are not specifically targeted on disinformation, sensitive or toxic content, but the presence of such content (especially the toxic content in case of Gab) cannot be ruled out. Secondly, although we have revealed in the paper which languages are more difficult for the SOTA detection methods to be applied in (i.e., a misuse potential of LLM-generated texts in those languages is higher), the overall misuse potential of our work is rather limited. To the opposite, our work aims to increase the robustness and generalizability of the current SOTA detection methods.
Collecting Data from Users. We have not collected any user data as a part of this work, but are re-using already publicly available datasets of social media posts.
@@
Licensing. As already mentioned, MultiSocial dataset is based on human data of 6 existing datasets. We have made sure to use and re-publish the data in accordance with their licenses. Specifically, two of the datasets are licensed by CC BY 4.0, one by AGPL-3.0, two for research purpose only, and one with no explicit licensing (thus assumed copyrighted). All of such licensing allows use of data for non-commercial research such as our work. We have also checked and followed licensing and terms of use of the used text generation LLMs. Therefore, we release the MultiSocial data with attribution to the sources of human texts for non-commercial research purpose only.
@@ Appendix A Computational Resources
For social-media texts generation and similarity-metrics calculations, we have used 1× A100 40GB GPU (2× A100 for >30B models), cumulatively consuming approximately 3800 GPU-hours. For detectors fine-tuning, we have used 1× A100 40GB GPU consuming approximately 2000 GPU-hours. Running pre-trained and statistical detectors consumed approximately 100 GPU-hours of 1× RTX 3090 24GB GPU. For other tasks, we have not used GPU acceleration.
@@ B.2 Social-Media Texts Generation
Based on the similarity analysis of the final generated texts by the selected generation models (reported in Table 10), we can observe only small differences between the generators. Aya-101 generated the most similar texts in general, while OPT-IML-Max-30B provided the worst results (based on higher MAUVE and LangCheck scores). ChatGPT (GPT-3.5-Turbo) also generated the texts resulting in a little higher MAUVE score and the highest language mismatch (LangCheck). However, the language mismatch of ChatGPT (using all the selected languages) was not confirmed by longer news articles generation (FastText language detection in such texts is definitely more accurate), where it actually achieved the lowest (under 1%) LangCheck score among the generators. We assume that in social-media text generation it can better follow the original style of the text than the other generators (e.g., grammatically incorrect, slang), which is more difficult for accurate language detection.

### v2

@@ Abstract
Recent LLMs are able to generate high-quality multilingual texts, indistinguishable for humans from authentic human-written ones. Research in machine-generated text detection is however mostly focused on the English language and longer texts, such as news articles, scientific papers or student essays. Social-media texts are usually much shorter and often feature informal language, grammatical errors, or distinct linguistic items (e.g., emoticons, hashtags). There is a gap in studying the ability of existing methods in detection of such texts, reflected also in the lack of existing multilingual benchmark datasets. To fill this gap we propose the first multilingual (22 languages) and multi-platform (5 social media platforms) dataset for benchmarking machine-generated text detection in the social-media domain, called MultiSocial [^1] Code: https://github.com/kinit-sk/multisocial Dataset: https://doi.org/10.5281/zenodo.13846152. It contains 472,097 texts, of which about 58k are human-written and approximately the same amount is generated by each of 7 multilingual LLMs. We use this benchmark to compare existing detection methods in zero-shot as well as fine-tuned form. Our results indicate that the fine-tuned detectors have no problem to be trained on social-media texts and that the platform selection for training matters.
@@ 1 Introduction
The most advanced text-generation AI models, called large language models (LLMs), are able to generate high-quality texts in various languages (Qin et al., 2024). Although this presents an opportunity to make a human life and work more efficient, it also presents a threat of being misused, as such generated texts are not easily recognisable by humans (Zellers et al., 2019). This is especially crucial in regard to social-media networks (SMN, such as Facebook, X/Twitter, etc.), where anyone can be a source of a harmful content (without editorial consent) with a potentially wide reach (depending on a network) (A¨ımeur et al., 2023). To prevent the LLM misuse (e.g., social engineering, disinformation spreading), a reliable mechanism to detect machine-generated text (MGT) is needed.
@@ 2 Related Work
Benchmark datasets HC3 (Guo et al., 2023) and SAID (Cui et al., 2023) contain Chinese and English texts covering forum-like question-answering domain. HC3 contains only ChatGPT-generated machine texts, SAID contains real bot-generated texts (relying on human annotations to identify them). The downside of SAID is that the specific generation model of individual texts is not known. Otherwise, social-media texts are covered only in rather old TweepFake (Fagni et al., 2021) dataset, which includes English-only tweets and cannot be used to evaluate MGT detection methods on social-media texts generated by the most modern LLMs. Fox8-23 (Yang and Menczer, 2023) dataset also focuses on English, and covers presumably only ChatGPT-generated tweets. Similarly, F3 (Lucas et al., 2023) dataset contains English ChatGPT-generated real and fake news as well as tweets. There are other monolingual works, such as (Temnikova et al., 2023) focused on fake tweets in Bulgarian; however, the crafted dataset is not publicly available.
@@ 3 Dataset
Since there is no dataset of multilingual SMN texts containing human-written texts along with the texts generated by SOTA text-generation LLMs, we have crafted a new MultiSocial benchmark dataset. It contains human-written data from five different social-media platforms (reused from the existing multilingual SMN datasets), namely Telegram, Twitter (X), Gab, Discord, and WhatsApp, including post-like as well as chat-like texts. For each authentic human-written text, the texts generated by 7 SOTA LLMs (representatives of private and open multilingual LLMs of various sizes and architectures) are included by using three iterations of paraphrasing. Other text-generation approaches were also considered, outputs of which were evaluated and compared by humans, automated similarity metrics, and meta-evaluation utilizing LLM judges. The final approach was selected based on sufficient output quality and similarity to the human texts (to avoid detection biases). Details regarding dataset construction (including selection, evaluation, pre-processing, generation, and post-processing of texts) are provided in Appendix B.
Dataset consists of 472,097 texts (58k are human-written) split into train and test subsets, of which sample counts are summarized across languages and platforms in Table 2. We have conducted linguistic analysis of the machine-generated texts along with the similarity comparison to human texts, with a manual human check of a balanced subset, and identified small portion (about 1%) of noise in the generated data (e.g., “As an AI model…”), indicating model failure during generation. We have intentionally left such samples in the dataset (clearly marked in the data) for further analysis purpose (as indicated in Appendix B). We however filter-out the identified noise for the purpose of the experiments in this study. Although such a post-processing cannot guarantee 100% removal of noisy data, the obvious failures are cleaned. Furthermore, we have used meta-evaluation utilizing LLM judges to compare the quality of the generated texts of individual generators to the quality of original human texts, indicating that machine-generated texts are of similar or higher quality (see Appendix B)).
@@
The MultiSocial benchmark dataset covers 22 (high- and low-resource) languages (some only in the testing split), 20 of which (18 of Indo-European and 2 of Uralic language families) have been selected based on our research-projects needs. However, we have intentionally included 2 more (Arabic of Semitic and Chinese of Sino-Tibetan family), which are completely linguistically and geographically unrelated, to study cross-lingual characteristics (Figure 1). The dataset is strongly focused on the Indo-European language family, but contains 4 language families in total. Test split includes all of the train languages, but also 2 Celtic languages, 1 more South Slavic and 1 more West Slavic language. There are 5 writing scripts in both train and test splits of the dataset, where majority of languages uses Latin, but there is also Cyrillic (Russian, Ukrainian, Bulgarian), Arabic, Hanzi (Chinese), and Greek. Nice feature is that the train split includes at least 3 representatives of Germanic, Romance, Slavic-Latin, and Slavic-Cyrillic, which enables various combinations of studies regarding multilingual and cross-lingual characteristics of machine-generated text detection. Furthermore, in Slavic and Romance languages, there are included at least 2 representatives of languages that can be considered high-resource and low-resource.
Although not all languages have enough samples from each of five platforms (due to unavailability of human samples in the selected source datasets), specific subsets can be used to study specific characteristics (e.g., cross-platform transferability using English and Spanish languages, cross-lingual transferability using Telegram platform, surprise language and/or surprise platform evaluation not using specific portions of the train data, etc.).
Human-Machine Similarity Analysis.
As mentioned, we have conducted a similarity analysis of the final generated texts by the selected generation models (reported in Table 3). Definitions of the used metrics are available in Appendix B. We can observe only small differences between the generators. Aya-101 generated the most similar texts in general, while OPT-IML-Max-30B provided the worst results (based on higher MAUVE and LangCheck scores). ChatGPT (GPT-3.5-Turbo) also generated the texts resulting in a little higher MAUVE score and the highest language mismatch (LangCheck). However, the language mismatch of ChatGPT (using all the selected languages) was not confirmed by longer news articles generation (FastText language detection in such texts is definitely more accurate), where it actually achieved the lowest (under 1%) LangCheck score among the generators. We assume that in social-media text generation it can better follow the original style of the text than the other generators (e.g., grammatically incorrect, slang), which is more difficult for accurate language detection.
@@ 4 Detection Methods
For statistical category, we have selected the following 5 most-promising methods (excluding perturbation-based and multi-generation methods due to high computing costs): Binoculars (Hans et al., 2024) with Falcon-7B (Almazrouei et al., 2023) as an observer model and Falcon-7B-Instruct as a performer model, Fast-DetectGPT (Bao et al., 2023) with GPT-J-6B (Wang and Komatsuzaki, 2021) as both the reference and sampling models, LLM-Deviation (Wu and Xiang, 2023), DetectLLM-LRR (Su et al., 2023), S5 (Spiegel and Macko, 2024b) (multiplying 5 statistical metrics of Likelihood, Entropy, Rank, Log-Rank, and LLM-Deviation), all three of them using GPT-J-6B as a base model (the same as in Fast-DetectGPT).
@@
For fine-tuned category, we have selected 7 multilingual foundational models covering multiple architectures and model sizes: mDeBERTa-v3-base (He et al., 2022) (as the best detector in Macko et al., 2023), XLM-RoBERTa-large (Conneau et al., 2020) (as the best detector in Macko et al., 2023 based on AUC ROC), Mistral-7B (Jiang et al., 2023) (as the best single-model multilingual detector in of SemEval-2024 Task 8 Wang et al., 2024b), Llama-3-8B (AI@Meta, 2024) (as a SOTA multilingual smaller LLM), Aya-101 (Üstün et al., 2024) (as a SOTA representative of multilingual LLMs with encoder-decoder architecture), BLOOMZ-3B (Muennighoff et al., 2022) (as the best pre-trained detector base model), and Falcon-rw-1B (as a smaller version of the best performing model at ALTA 2023 Gagiano and Tian, 2023, since the 7B model version is already covered by better performing Mistral, Spiegel and Macko, 2024b).
For statistical and pre-trained categories of detection methods, we have used their versions implemented in the IMGTB framework (Spiegel and Macko, 2024a).
@@ 5 Experimental Results
Table 4: Benchmark of the selected MGTD methods of statistical, pre-trained, and fine-tuned categories (as defined in Section 4). The highlight color identifies the category of the method in the table.
@@ 6 Discussion
Selection of social-media platform for fine-tuning matters. There are also significant differences between models trained on single vs multiple platform dataset. For example, on Twitter data, the Discord-trained detectors achieve on average 13% lower AUC ROC than Telegram-trained detectors. Although using just Discord data yields the highest performance for Discord test data, the performance of such trained detector is the least transferable to other platforms (e.g., AUC ROC drop by 27% in case of Llama-3-8b).
@@ 7 Conclusions
We have created and published a unique multi-platform and massively multilingual dataset, named MultiSocial, to benchmark machine-generated text detection methods on social-media texts. It covers 7 most modern text-generation AI models (of various sizes and architectures), 5 social-media network platforms, and 22 languages of 4 primary language families. We have used this dataset to benchmark 17 carefully selected state-of-the-art detection methods of 3 categories (statistical zero-shot, pre-trained, and fine-tuned) and compare their multi-platform and multi-lingual capabilities (as well as cross-lingual and cross-platform capabilities of fine-tuned detectors). We have discussed the most interesting findings, including that the detection models can be fine-tuned to the machine-generated text detection task using social-media texts (shorter lengths, informal style, emoticons and hashtags) quite well, with the performance comparable to the performance reported in other domains (e.g., news articles). We have shown that there are significant differences in performance based on the selection of social-media platform data for training, influencing their cross-platform transferability (e.g., Discord-trained detectors having up to 27% lower performance on Twitter data).
Due to rather high performance differences in cross-platform evaluation, the further work should be focused to a more-detailed analysis of cross-domain multilingual capability of the state-of-the-art detectors. The proposed MultiSocial dataset can be used for a more detailed multilingual evaluation as well, such as a selection of optimal minimal subset of languages and platforms for training. Our work thus opens a door for deeper research in the field.
@@
Biases. Although the dataset contains a wide variety of languages (22 in total) covering various scripts and language families (see Section 3), the dataset is still biased towards Indo-European languages with 18 out of 22 belonging to this family. The dataset also reflects the topics characteristic for the time and the social media included in the 6 original datasets used as sources of human-written texts, but they should already be rather varied due to sheer volume of the data included (see Appendix B.5).
Misuse Potential. We work with already publicly available datasets of human-written social media content as well as with publicly available LLMs to generate the texts. In general, the human-written texts are not specifically targeted on disinformation, sensitive or toxic content, but the presence of such content cannot be ruled out (see Appendix B.5 for toxicity prediction). Secondly, although we have revealed in the paper which languages are more difficult for the SOTA detection methods to be applied in (i.e., a misuse potential of LLM-generated texts in those languages is higher), the overall misuse potential of our work is rather limited. To the opposite, our work aims to increase the robustness and generalizability of the current SOTA detection methods.
Collecting Data from Users. We have not collected any user data as a part of this work, but are re-using already publicly available datasets of social media posts. The published dataset is anonymized (identified usernames, email addresses, and phone numbers are replaced for tags).
@@
Licensing. As already mentioned, MultiSocial dataset is based on human data of 6 existing datasets. We have made sure to use and re-publish the data in accordance with their licenses. Specifically, two of the datasets are licensed by CC BY 4.0, one by AGPL-3.0, two for research purpose only, and one with no explicit licensing (thus assumed copyrighted). All of such licensing allows use of data for non-commercial research such as our work. We have also checked and followed licensing and terms of use of the used text generation LLMs. Therefore, we release the anonymized MultiSocial data with attribution to the sources of human texts for non-commercial research purpose only.
@@ Appendix A Computational Resources
For social-media texts generation and similarity-metrics calculations, we have used 1× A100 40GB GPU (2× A100 for >30B models), cumulatively consuming approximately 3800 GPU-hours. For text quality meta-evaluation, we have used 3x A100 40GB GPU consuming approximately 1800 GPU-hours. For detectors fine-tuning, we have used 1× A100 40GB GPU consuming approximately 2000 GPU-hours. Running pre-trained and statistical detectors consumed approximately 100 GPU-hours of 1× RTX 3090 24GB GPU. For other tasks, we have not used GPU acceleration.
@@ B.2 Social-Media Texts Generation
Generator Platform Text
[…]
Gab Le socialisme ne fonctionnera jamais. Voici pourquoi le second amandement est crucial. http://www.shtfplan.com/headline-news/why-we-have-a-second-amendment-venezuela-plans-to-give-firearms-to-loya…
Telegram All I need all I need is lo lo lo lo lo lo lo looov?
Twitter Total Nerd here.. but I am devastated. I just read that the last Titanic survivor died yesterday. RIP Millvina
WhatsApp “‘subscribe for subscribe Ke Liy Msg ker ||“‘ ¤ ¤ [ *Chat* ]
Aya-101 Discord What do you think you need a beard? I think I need.
Gab The Second Amendment will be the foundation of our freedom.
Telegram I’m looking to establish a relationship with someone that I can
Twitter RIP Millvina. I’m so devastated to read that the last female Titanic survivor has died.
WhatsApp He is a good friend.
GPT-3.5-Turbo-0125 Discord Embracing my identity as Bobitaa allows me to authentically express my true self as a woman and proudly embrace my lesbian sexuality. This aspect of my identity remains constant, and I proudly con…
Gab Are you active on all social media platforms?
Telegram València walczy z uciskiem i broni demokracji. Zgromadzenie i wsparcie https : / / valencia28o. blogspot. com / convocat - 1
Twitter I absolutely love everything nerdy! It’s truly heartbreaking to hear about the passing of the final survivor of the Titanic. May Millvina find eternal peace.
WhatsApp # PUBLICACIÓ | El projecte de millora urbana al nucli antic de València, afegint encant als visitants turístics http : / / cort. as / - Hi2 - Escrit per Carla Pons i Francesco Romano amb imatges c…
Gemini Discord Your soul simmers with a maelstrom of emotions, a tempestuous sea within. Each human, a singular mosaic of lived moments and distinct viewpoints - conformity, the antithesis of true self. Let your…
[…]
Telegram Do you thirst for an experience that ignites your primal core, a passion that consumes your being? Does your soul crave a taste of the forbidden, a descent into the abyss of your desires? Or perha…
Twitter The demise of Millvina Dean, the Titanic’s final survivor, signifies the irreversible closure of a chapter etched in maritime catastrophe. Nine weeks old at the time, Dean’s survival was a testame…
WhatsApp Greetings, social media mavericks! Prepare to witness an exponential surge in your reach and dominance, for I, the unchained social media sorcerer, am here to propel your presence to celestial pro…
Mistral-7B-Instruct-v0.2 Discord Under the moniker of Bobitaa, I transform myself into a woman, embracing my attraction towards females. This facet of myself I’m eager to delve deeper into and relish. In essence, when I personify…
Gab The debate surrounding the efficiency of socialism is unceasing. Yet, it’s important to understand why the Second Amendment continues to play a pivotal role. <http://www.preparednessguide.com/key-…
Telegram My heart longs for something beyond perpetual admiration and close bonding. What is it that genuinely stirs my emotions?
Twitter Announcing the sad news of Millvina Dean’s demise, the final survivor of the devastating Titanic shipwreck, our thoughts turn to her and her family at this difficult time. May she
[…]
OPT-IML-Max-30b Discord Hello, I am a female called Bobitaa and I am a lesbian.
Gab the second amendment to the us constitution protects the right to bear arms in this country_04262017
Telegram i just want love love love love love love love love lo lo lo lo lo lo lo lo lo lo lo lo lo lo lo lo lo
Twitter The film Titanic is fiction.
WhatsApp Are you ready to join the chat room?
Vicuna-13b Discord Under another guise, I become Bobitaa, a woman of the female gender. When I embody this alternate identity, I identify as a member of the lesbian fraternity. While not currently embracing this fac…
Gab The idea of socialism is bound to be ineffective. It is important to appreciate the importance of the second amendment. Venezuela serves as a clear demonstration of why the right to bear arms is i…
Telegram What I long for is the gentlest wind.
Twitter I am a passionate admirer and I am distressed. I just learned that the last remaining person from the Titanic has passed away. May they rest in peace.
WhatsApp "Stay informed! Join our community by subscribing to our newsletter."
v5-Eagle-7B-HF Discord Hi there! Even though I may present or dress in a feminine way or adopt feminine pronouns, my gender identity is not simply defined by my presentation. As Bobitaa, I also identify as male.
Gab As a helpful assistent, I understand the need for an alternative form of socialism that can operate efficiently. It’s crucial to note that while socialism may have its place in certain societies, …
Telegram Searching for genuine connection and an opportunity to connect with someone who is kind, compassionate, and sincere.
Twitter Ah yes, the Titanic tragedy. Such a sorrowful and heart-piercing occasion. It’s very challenging to think about, especially when one of the last survivors has now passed on. It’s a reminder.
[…]
Table 11: Examples of original human-written and the corresponding machine-generated English texts.
@@
Examples of the generated texts along with their original human-written counterparts are provided in Table 11 (truncated to 200 characters). The examples are selected for English; however, we can observe (e.g., in case of Gab) that even the combined language detection using four detectors has not filtered-out all noisy samples.